### v2

# NativQA Framework: Enabling LLMs with Native, Local, and Everyday Knowledge

Abstract
[…]
NativQA Framework:
Enabling LLMs with Native, Local, and Everyday Knowledge
Firoj Alam,1 Md Arid Hasan,2 Sahinur Rahman Laskar,3 Mucahid Kutlu,4 Kareem Darwish,1 Shammur Absar Chowdhury1 1Qatar Computing Research Institute, Qatar, 2University of New Brunswick, Canada, 3UPES, India, 4Qatar University, Qatar fialam@hbku.edu.qa, arid.hasan@unb.ca https://gitlab.com/nativqa/nativqa-framework
[…]
NativQA Framework:
Enabling LLMs with Native, Local, and Everyday Knowledge
Firoj Alam,1 Md Arid Hasan,2 Sahinur Rahman Laskar,3 Mucahid Kutlu,4 Kareem Darwish,1 Shammur Absar Chowdhury1 1Qatar Computing Research Institute, Qatar, 2University of New Brunswick, Canada, 3UPES, India, 4Qatar University, Qatar fialam@hbku.edu.qa, arid.hasan@unb.ca https://gitlab.com/nativqa/nativqa-framework
1 Introduction
(a) Language and region coverage.
(b) Topics.
Figure 1: The NativQA framework has been used to collect QA pairs across 39 locations in 24 countries, covering 7 languages. (a) The plot shows city-level data distribution grouped by language—inner circles denote languages, outer wedges represent cities (with country names), sized by sample count. (b) Topics evaluated with the corresponding datasets are also shown.
Large Language Models (LLMs) have exhibited cultural bias toward certain cultures, largely due to the predominance of training data in high-resource languages and the limited digital representation of low-resource languages Li et al. (2024a); Pawar et al. (2024); Durmus et al. (2024). Given that most LLMs are trained primarily on English or translated data, which often fails to capture how people naturally speak, ask questions, and express their needs in their native languages and dialects. Since the emergence of LLMs, there has been growing concern about how to develop and benchmark models that perform equitably across diverse languages and cultures Li et al. (2024b); Myung et al. (2024); Kim et al. (2024); Pawar et al. (2024). In addition to cultural representation, [^1] Cultural ability includes understanding the diversity of cultural elements across different cultures Pawar et al. (2024). In the context of cultural QA, it can be defined as question answering that reflects the types of questions likely to be asked and answered by people of a particular culture, either in their native or second language Arora et al. (2024). there is also concern about whether LLMs can effectively respond to everyday, long-form, and complex queries that require paragraph-length answers Arora et al. (2024); Myung et al. (2024); Hasan et al. (2024). These concerns have driven efforts to develop multilingual Üstün et al. (2024); Touvron et al. (2023); Team et al. (2024) and language-specific Team et al. (2025); Nahin et al. (2025); Nguyen et al. (2024) closed and open-source LLMs. To benchmark and fine-tune models with region- and culture-specific knowledge, there have been efforts to develop language- and culture-specific resources, along with benchmarked and fine-tuned models Chiu et al. (2024); Shi et al. (2024); Li et al. (2024a). Methodologies involved in these efforts include collecting QA pairs from community question-answering forums Shi et al. (2024), hiring native speakers to write questions and answers Chiu et al. (2024), translating content from English into local languages Myung et al. (2024), and using LLMs to generate QA pairs Putri et al. (2024). A common finding across studies is that while most LLMs perform reasonably well on high-resource languages, they tend to struggle with mid- and low-resource languages Myung et al. (2024). Several studies have demonstrated that fine-tuning with culture-specific knowledge can improve model performance, even in smaller or quantized models Li et al. (2024a, b); Hasan et al. (2024). These findings suggest that achieving inclusivity in LLMs requires datasets that reflect diverse regions, cultures, and content generated by native users. One of the major challenges in creating linguistically and culturally rich datasets is the significant manual effort involved. As LLMs continue to advance, the need for scalable methods to build these resources becomes more critical. However, there is currently a lack of tools designed to support the efficient development of such datasets. To bridge this gap, we propose NativQA framework–a system that enables communities to easily build culture-specific QA datasets for enriching and benchmarking LLMs. We evaluated NativQA framework for 39 locations, from 24 countries covering 7 languages and 18 topics, which resulted over 300K data points as shown in Figure 1. To summarize, the contributions of our study are as follows.
  • We propose NativQA framework, for creating culture- and region-specific QA datasets that promote inclusivity in LLMs by enabling community-driven development of diverse, culturally aligned benchmarks.
  • The NativQA framework offers several key features: (i) seamless integration of both user- and LLM-generated queries to collect QA pairs; (ii) location-agnostic – users can collect data from any region; (iii) support for multiple search engines; (iv) image search capability; (v) automated removal of duplicate QA pairs; (vi) domain reliability checks; and (vii) an efficient caching mechanism to reduce redundant API calls, leading to cost savings and mitigating time-out issues.
  • We conducted an extensive evaluation to assess its feasibility for broader adoption and demonstrated that it can be effectively used across different locations, cultures, and languages.
  • We have made the NativQA framework publicly available for the community.
2 NativQA framework
In Figure 2, we present the NativQA framework consisting of three inter-connected modules such as (i) Query Collection, (ii) QA Collection, and (iii) QA Validation. The top part of the figure shows the entire process while the bottom part shows an example from query collection to QA validation. In the below sections we discuss them in details.
Figure 2: NativQA framework, demonstrating the entire pipeline, from query collection to the final dataset development process.
2.1 Query Collection (QC) Module
The objective of this module is to collect open-ended queries focused on various predetermined topics derived from common knowledge in everyday communication. The set of topics should be established prior to beginning the query collection process, as this facilitates the gathering of topic-specific queries. As shown in Figure 1, the framework has been evaluated using a predetermined set of 18 topics that are culture- or region-dependent (e.g., Events, Literature, etc.). However, the topics can be anything and can be task specific. The query collection process can be carried out using three different approaches:
  • Manual query collection: Involves asking annotators to write topic-specific queries they might pose when seeking everyday information.
  • Template-based generation: Defines templates for different topics with a placeholder for the location. For example, “main cultural festivals in [LOCATION]”, where LOCATION can be a city, country, or region name.
  • LLM-based generation: Uses LLMs to generate queries.
The manual query collection process requires recruiting location- or region-specific annotators. Moreover, like any other annotation task, it is time-consuming. For example, in Hasan et al. (2024), country-specific annotators were hired to manually write topic specific queries. In contrast, the latter approaches are more cost- and time-effective. The template-based approach involves defining location-agnostic templates, which can be seamlessly adapted for different locations. The LLM-based approach requires prompt engineering to generate appropriate queries. We use the term seed query ($Q_{0}$) to refer to the queries generated by all approaches. These queries are then used in later steps by the QA collection module to gather QA pairs and additional queries. Filtering: Given that queries can have exact and/or near duplicates, it is important to remove such duplicates as part of the filtering process. The filtering module handles this by using exact matching for exact duplicates and a similarity-based approach for near duplicates. These steps results in the final set of queries.
2.2 QA Collection (QAC) Module
The next step is to collect QA pairs using a search engine, e.g., Google. Most search engines support features like “People also ask” or “Related queries,” which list questions asked by real users that are potentially relevant to the initial seed query, as shown in Figure A1. Moreover, these questions are accompanied by answers extracted by the search engine, along with links to the sources of the answers. The QA collection module implements Algorithm 1 (outlined in Appendix). It takes the seed queries $Q_{0}$ and the number of iterations $N_{\text{iter}}$ as input. For each iteration $i\in N_{\text{iter}}$, it collects QA pairs $P_{QA}^{i}$ and related queries $S^{i}_{\text{rel}}$ for each query $q\in Q$, then passes them to the filtering module and updates the current query set $Q$. This process is repeated for all iterations to obtain the final QA set, $S_{QA}$, for the enriched queries $Q$. The idea of the filtering module is to remove duplicates that may arise in different iterations. The QA collection module leverages search engine APIs to collect QA pairs, specifying location and language parameters in the API call. These parameters define the search criteria. Location parameter is especially important, as the focus is on collecting QA pairs from specific regions. The module also stores relevant metadata (e.g., answer source URL) for later use.
2.3 QA Validation (QAV) Module
The next step of the framework is to validate the QA pairs. It is important for several reasons: (i) the retrieved related queries might be less relevant, (ii) answers might be incomplete, (iii) the source of the answer can be less reliable. For the QAV the framework integrates the below two approaches.
1. Domain Reliability Checking (DRC).
The answers collected by NativQA framework include a link to the source web page from which the answer was extracted. Therefore, the aim of the DRC is to retain QA pairs based on the reliable web domain where the answer appears. We hypothesize that answers from pages hosted on reliable domains are more likely to be trustworthy and can be directly included in the dataset. The framework currently supports two options for DRC:
  • Based on the domains of the QA pairs listed in $S_{QA}$, annotators manually check and label them as very reliable, partially reliable, not sure, or completely unreliable. For this task, the annotation guidelines proposed in Hasan et al. (2024) can be adopted.
  • In the current implementation, NativQA framework provides $2,080$ source URLs obtained from the MultiNativQA dataset Hasan et al. (2024), where annotators identified them as very reliable. These domain URLs used to verify the reliability of the answers.
This approach is more practical and scalable for developing large-scale QA pairs, as it reduces the manual effort required to obtain hand-curated answers. However, due to the semi-supervised nature of the approach, it has certain limitations. Some domains are authentic (e.g., Google, Facebook); however, the content they host might not be reliable due to the questionable user-generated content.
QA Annotation (QAA).
To increase more reliability of the curated QA pairs, different steps can be taken. Currently NativQA framework supports two different approaches:
  • Manual annotation: This approach involves manually reviewing and editing the answers. It includes three types of annotations. (i) Question validation: Human annotators assess the quality of the questions by assigning one of the following labels: good or bad. A “good question” is defined as a fact-seeking question that can be answered with an entity or an explanation. In contrast, a “bad question” is either ambiguous, incomprehensible, based on a clear false presupposition, opinion-seeking, or does not seek factual information. (ii) Answer editing: If a given answer does not fully address the question, or includes extra or incorrect information, the annotator must edit the answer using content from the source web page to ensure completeness and accuracy. Annotators can only use the provided source web pages to maintain the scope and the reliability of the answers. (iii) Location relevance: This task involves determining whether the question is relevant to the specified [LOCATION].
  • LLM-based annotation: While manual annotation by humans is more reliable, it is also time-consuming and costly. To reduce such effort, NativQA framework supports the use of LLM-based approaches as models like GPT-4 have demonstrated effectiveness in improving task-specific model performance Li et al. (2023, 2024c). In the LLM-based approach, a concise version of the annotation guidelines Hasan et al. (2024) and task definitions is adapted into the prompt to check the question, edit the answer, and identify location relevance.
2.4 Executing the Framework
Once the seed queries are curated, running NativQA framework is straightforward – executed with a single command and customizable parameters. It reads the seed queries from the specified input file (--input_file <$file$>) and, for each query, calls the search API with additional parameters such as --country_code <$c$>, --location <$l$>, and the number of iterations $n$ (--limit <$n$>). An example command is provided below. The NativQA framework is equipped with other scripts (e.g., domain reliability checking), which are detailed as a part of the framework’s documentation.
\$python -m nativqa --engine google --search_type text --input_file <data/test_query.csv> --country_code qa --location <"Doha, Qatar"> --env <envs/api_key.env> --n_iter n
3 Evaluation of NativQA framework
Evaluated in multiple studies. The NativQA framework used in Hasan et al. (2024) to collect data from 9 locations in 7 languages, where seed queries and QA validation performed manually. This effort resulted in approximately 64K annotated QA pairs. Mousi et al. (2025) used the framework for Arabic cultural and dialectal QA data collection to benchmark the cultural and dialectal capabilities of LLMs. Evaluation through additional data collection. We further evaluated the framework for additional 30 locations by creating seed query templates. We developed Arabic and English seed query templates from the seed queries that were used to collect the MultiNativQA dataset Hasan et al. (2024). We selected seed queries that contain mentions of locations in the MultiNativQA, which were then used with the [LOCATION] template (e.g., women’s wedding attire in [LOCATION]). The [LOCATION] template was subsequently replaced with other locations to create seed queries for different regions. We manually validated the seed queries generated from the template and removed those that were specific to a particular location (e.g., 1971 liberation war in Bangladesh/[LOCATION]). For Arabic, we used only the seed queries from the template that were employed to collect QA data for Qatar, while the English queries were adopted from the manual seed queries for both Bangladesh and New York. In Appendix C, we discuss the details of the data distribution (over 300K) that have been collected using the NativQA framework. Capitalizing LLMs. In the NativQA framework, we used LLMs in two stages: (i) query generation and (ii) LLM-based QA annotation. The use of LLMs for such tasks involves prompting them to produce the desired outcomes. We will release the prompts and scripts as part of the framework. LLM-based Query generation. For query generation, we designed prompts specifying the location and topic, and used LLMs to generate queries. To assess the quality of the generated queries, we manually reviewed a random sample and found that they were acceptable to human annotators (as they matched human-generated queries) and reflected the specified location and topic. LLM-based QA annotation. For LLM-based QA annotation, a concise version of the annotation guidelines and task definitions is adapted to the prompt to validate the question, edit the answer, and identify the location relevance. To manually assess the quality of the annotated QA pairs, we choose two different approaches: (i) QA validation by selecting the best answer, and (ii) QA annotation for LLM-based QA annotation.
  • Manual QA validation - selecting best answer. For QA validation for LLM-based QA annotation, we manually evaluated the output of LLMs on 500 QA pairs collected from Egypt. We selected this set due to the availability of annotators from Egypt. The goal was to assess whether human annotators preferred the LLM-edited answers or the ones retrieved from a search engine. We provided clear instructions to the annotators to select the better answer based on accuracy, clarity and readability, as well as completeness and relevance (see in Appendix D.1). Annotators were given three options: (i) Answer 1 (original answer), (ii) LLM-edited answer, and (iii) neither. We recruited two annotators who are native Arabic speakers residing in Egypt. They were compensated at a standard rate managed by a third-party company. We computed inter-annotator agreement using observed agreement and Gwet’s AC1 Gwet (2008), and Cohen’s Kappa. The observed agreement and Gwet’s AC1 were 0.842 and 0.806, respectively, indicating strong overall alignment. The high observed agreement suggests that using LLMs to edit answers may be a promising direction.
  • Manual QA validation - answer scoring. This approach involves manually checking the quality of LLM-edited answers and scoring them. We used a 5-point Likert scale for various evaluation metrics, including clarity, faithfulness, informativeness, and plausibility. These metrics are selected from various related studies on the evaluation of natural language explanations Huang et al. (2024); Zavolokina et al. (2024). We provide detailed annotation guidelines for human evaluators in Appendix D.2. Due to the availability of native human annotators, the evaluation was carried out for two locations, such as Sudan and Yemen, on $\sim$350 examples from each test set. Each LLM-edited answer was annotated by three and two human evaluators for Sudan and Yemen, respectively. We presented the average Likert-scale scores for all the evaluation metrics in Table 1. Furthermore, we measured annotation agreement on ordinal scales using the agreement index $r^{*}_{wg(j)}$ James et al. (1984), which evaluates how the observed rating variance compares to the maximum variance expected under total disagreement. The values over 0.85 for Sudan and 0.86 for Yemen indicate a strong agreement O’Neill (2017), as shown in Table 1. In addition, we analyze the annotations and find that 100% answers were not edited by annotators for the Sudan data. For Yemen, the percentage of not edited answer was 92%. Both human annotation results and analysis strongly suggest that LLMs are highly capable of QA annotation in the NativQA framework.
Location Clarity Faithfulness Informativeness Plausibility
Average Likert score
Khartoum, Sudan 4.41 4.42 4.26 4.29
Sanaa, Yemen 4.05 4.05 3.98 4.06
Agreement index $r^{*}_{wg(j)}$
Khartoum, Sudan 0.85 0.85 0.85 0.85
Sanaa, Yemen 0.88 0.86 0.89 0.88
Table 1: Average Likert scale and agreement scores for each human evaluation metric for LLM-edited QA.
For both query generation and LLM-based QA annotation, we used GPT-4o (version: 2024-11-20). However, any models can be used for such tasks.
4 Features of NativQA framework
NativQA framework features a generic framework that is agnostic to both location and language. It enables a scalable regional and language data collection process while offering simplicity in implementation and flexibility in customization. Modularity. The NativQA framework, as shown in Figure 2, follows loosely coupled design principles, effectively separating the seed query collection, QA curation, and QA validation components. Their interactions ensure a modular and flexible architecture. Generality. The framework is designed to offer generality, allowing effortless customization of location, language, preferred search engines, and modalities. It supports both text- and image-based search functionalities. Data, along with metadata, are stored in JSON and JSONL formats. Overall, the framework provides a high level of generality and can be readily applied and adapted to a wide range of use cases and research scenarios. Caching. One of the significant challenges when accessing APIs is managing timeouts and other failure issues. This often requires re-running experiments involving API calls, which not only demands additional effort but also increases costs. To address this problem, we have developed a caching mechanism that allows users to bypass API calls for queries that have already been successfully processed. Specifically, all intermediate outputs are saved when processing a data sample. During a re-run, samples with cached model responses do not access the API. Open-source. We made NativQA framework accessible to the community by releasing it as open-source. This will also enable the continued development of the framework within the community.
5 Related Work
Local and Cultural Knowledge in LLMs.
LLMs excel on standard NLP benchmarks Achiam et al. (2023); Touvron et al. (2023); Bubeck et al. (2023); Bang et al. (2023); Ahuja et al. (2023); Hendy et al. (2023); Abdelali et al. (2024), yet often struggle with tasks requiring culturally grounded knowledge, especially in low-resource languages and dialects Pawar et al. (2024). This gap highlights the need for more inclusive models. Recent studies demonstrate the importance of enabling LLMs to understand and generate content reflective of diverse linguistic forms, social norms, and local experiences Li et al. (2024b, a); Shi et al. (2024). A culturally inclusive model should be capable of understanding topics like healthcare, education, or cuisine within local contexts. However, current LLMs frequently fall short in capturing region-specific expressions and indigenous knowledge Myung et al. (2024); Chiu et al. (2024).
Cultural, Information-Seeking, and Multilingual Resources.
Focusing on cultural aspects, there have been efforts to develop datasets that benchmark and fine-tune LLMs for diverse languages and regions Arora et al. (2024); Shi et al. (2024); Li et al. (2024b). In addition to cultural knowledge, there have also been initiatives to build resources that reflect everyday information-seeking behavior Myung et al. (2024). Efforts targeting everyday information-seeking queries include the development of mono- and multilingual resources by collecting data from diverse sources such as Wikipedia Yang et al. (2018); Kwiatkowski et al. (2019), Google Search QA Khashabi et al. (2021), Reddit forums Fan et al. (2019), and question–answer pairs written directly by native speakers Clark et al. (2020). Other approaches include combining native and machine-translated content Arora et al. (2024), and generating QA using LLMs Putri et al. (2024); Li et al. (2024b).
Motivation for NativQA framework.
Prior work has advanced the inclusion of native language questions and culturally specific content. Datasets like TyDi QA Clark et al. (2020), BLEnD Myung et al. (2024), and CaLMQA Arora et al. (2024) broaden cultural coverage, while search engine-based datasets such as NQ Kwiatkowski et al. (2019), MS MARCO Nguyen et al. (2016), and GooQA Khashabi et al. (2021) offer scalability and real-world relevance, albeit primarily in dominant languages. Despite these advances, a unified framework addressing multilinguality, location specificity, and scalability remains missing. NativQA framework bridges this gap by enabling large-scale data collection from real-world search queries with culturally grounded, native phrasing in underrepresented languages.
6 Conclusions and Future Work
In this study, we propose the NativQA framework, which a scalable, language-independent framework to build culturally and regionally aligned QA datasets in native languages. It utilizes search engines capabilities to collect QA based on seed queries that native users ask. Based on the seed queries and target search locale one can collect a large number of QA pairs relatively quickly. The framework has already been used in published studies. We additionally evaluated its utility for another 30 locations from different countries and languages. Its functionality for multi-modalities can enable further research in this area and caching mechanism will help in reducing the search engines API cost. In future we aim to extend NativQA framework for other modalities such as audio/video and evaluate for more under-representative locations.
7 Limitations
The proposed NativQA framework enables the development of datasets with cultural and native information for everyday knowledge. The NativQA framework currently uses human-in-the-loop processes in different phases such as seed query creation and manual revision of QA pairs. Though it also supports LLMs for seed query generation and QA validation, however, it is important to have a human-in-the-loop setup to verify the LLMs outputs. The framework does not include a built-in annotation platform for manual QA verification. However, any annotation platform (e.g., Pybossa [^2] https://github.com/Scifabric/pybossa) can be adapted for this purpose.
Ethics Statement
The proposed NativQA framework does not involve collecting any personally identifiable information. Therefore, we do not foresee any issues that may lead to potential risks.
Appendix
Appendix A QA from Search Engines
In Figure A1, we present a screenshot of a Google search results page for the query “visit Baladna Farm in Qatar.” The “People also ask” section contains related questions, which we used to collect question-answer (QA) pairs. These questions were subsequently used as queries to collect additional QA pairs.
Figure A1: QA list in response to a query obtained from Google.
Appendix B Algorithm for QA Collection
Algorithm 1: Collecting QA Pairs from Search Queries. For QA collection we report Algorithm 1. It outlines a bootstrapped approach for collecting question-answer (QA) pairs using an initial set of seed queries $Q_{0}$. The process is iterative, aiming to expand the query set and extract more QA pairs at each step. The algorithm takes as input a predefined set of seed queries $Q_{0}$ and the number of iterations $N_{\text{iter}}$. At each iteration, for every query in the current set $S\varrho$, we retrieve QA pairs and related queries using the functions ExtractQA(*) and ExtractRelatedQueries(*), respectively. ExtractQA(*) returns a set of question-answer pairs $P_{QA}^{i}$ with corresponding attribution labels $L$, while ExtractRelatedQueries(*) retrieves additional queries that are semantically or topically related to the input query. Duplicate QA pairs are removed through the DeDuplication(*) function to ensure the uniqueness of the collected data. The new QA pairs are added to the global set $S_{QA}$, and the related queries are incorporated into the query pool $S\varrho$ for use in subsequent iterations. This approach enables iterative enrichment of both the query pool and the QA dataset, facilitating broader coverage of topics while maintaining data quality through deduplication.
Algorithm 1 Collecting QA pairs using seed queries $\varrho_{0}$. $P_{QA}^{i}$: QA pair, $S\varrho_{rel}^{i}$: related queries. ExtractQA(*) and ExtractRelatedQueries (*) are functions that return questions, $Q$-answers, $A$ pairs with attribution $L$, and related queries, respectively, which are obtained from the search engine for a given query, $q$. DeDuplication (*) removes any duplicate entries from the set to ensure uniqueness.
1:Input:
2:      Seed queries: $\varrho_{0}=\{\hat{\varrho_{1}},\hat{\varrho_{2}},\ldots,\hat{\varrho_{m}}\}$
3:      Number of iterations: $N_{iter}$
4:Output:
5:      Set of QA pairs: $S_{QA}$
6:      Set of enriched queries: $S\varrho$
7:$S_{QA}\leftarrow\emptyset$
8:$S\varrho\leftarrow\varrho_{0}$
9:for $i$ from 1 to $N_{iter}$ do
10:     $P_{QA}^{i}\leftarrow\emptyset$
11:     $S\varrho_{rel}^{i}\leftarrow\emptyset$
12:     for $q\in S\varrho$ do
13:         $(Q^{q},A^{q},L^{q})\leftarrow\text{ExtractQA}(q)$
14:         $P_{QA}^{i}\leftarrow P_{QA}^{i}\cup\{(q^{\prime},a^{\prime},l^{\prime})\mid q^{\prime}\in Q^{q},a^{\prime}\in A^{q},l^{\prime}\in L^{q}\}$
15:         $S\varrho_{rel}^{i}\leftarrow S\varrho_{rel}^{i}\cup\text{ExtractRelatedQueries}(q)$
16:     end for
17:     $P_{QA}^{i}\leftarrow\text{DeDuplication}(P_{QA}^{i})$
18:     $S_{QA}\leftarrow S_{QA}\cup P_{QA}^{i}$
19:     $S\varrho\leftarrow S\varrho\cup S\varrho_{rel}^{i}$
20:end for
21:return $S_{QA},S\varrho$
Appendix C Distribution of Additional Dataset
Language Location Train Dev Test Total
Arabic Abu Dhabi, UAE 4,896 692 1,406 6,994
Algiers, Algeria 15,198 2,149 4,364 21,711
Baghdad, Iraq 17,910 2,533 5,143 25,586
Beirut, Lebanon 17,748 2,510 5,096 25,354
Cairo, Egypt 8,631 1,221 2,478 12,330
Damascus, Syria 7,902 1,117 2,269 11,288
Doha, Qatar 3,649 492 988 5,129
Gaza, Palestinian 3,975 562 1,142 5,679
Khartoum, Sudan 3303 467 948 4,718
Kuwait City, Kuwait 21,676 3,066 6,224 30,966
Manama, Bahrain 17,599 2,489 5,053 25,141
Muscat, Oman 3,479 492 999 4,970
Nouakchott, Mauritania 13,520 1,912 3,882 19,314
Rabat, Morocco 17,341 2,453 4,979 24,773
Riyadh, Saudi Arabia 4,150 587 1,191 5,928
Sanaa, Yemen 3,373 477 968 4,818
Tripoli, Libya 3,300 467 947 4,714
Tunis, Tunisia 10,352 1,464 2,972 14,788
Assamese Assam, India 1,131 157 545 1,833
Bangla Dhaka, Bangladesh 7,018 953 1,521 9,492
Kolkata, India 6,891 930 2,146 9,967
English California, USA 1,065 151 306 1,522
Dhaka, Bangladesh 4,761 656 1,113 6,530
Doha, Qatar 8,212 1,164 2,322 11,698
Florida, USA 777 110 223 1,110
Georgia, USA 816 115 234 1,165
Hawaii, USA 788 112 226 1,126
Illinois, USA 877 124 252 1,253
Massachusetts, USA 842 119 242 1,203
Michigan, USA 796 113 228 1,137
New York, USA 4,518 623 1,313 6,454
North Carolina, USA 809 115 232 1,156
Ohio, USA 862 122 247 1,231
Ontario, Canada 743 105 213 1,061
Pennsylvania, USA 786 111 226 1,123
Quebec, Canada 720 102 207 1,029
Texas, USA 847 120 243 1,210
Washington, USA 874 124 251 1,249
Hindi Delhi, India 9,288 1,286 2,745 13,319
Nepali Kathmandu, Nepal 561 561
Turkish Istanbul, Turkey 3,527 483 1,218 5,228
Total 234,950 33,045 67,863 335,858
Table A1: Statistics of the data collected using NativQA Framework. – Only testing split due to limited dataset size.
In Figure A2, we show the distribution of the dataset across major cities and languages. The QA dataset covers seven languages—Arabic, English, Hindi, Bangla, Turkish, Assamese, and Nepali—capturing a rich linguistic landscape spanning 39 global cities. Arabic QA pairs come from a variety of MENA countries, including Egypt, Iraq, Lebanon, and Qatar. English content is drawn from both Western countries (e.g., the USA and Canada) and multilingual regions such as Qatar and Bangladesh. South Asian languages like Hindi, Bangla, Assamese, and Nepali are represented through key urban centers including Delhi, Dhaka, Assam, Kolkata, and Kathmandu. Moreover, we present the details of the collected data using NativQA framework in Table A1. We split the data into training (70%), development (10%), and test (20%) sets.
Figure A2: Distribution of the dataset collected using NativQA framework across different languages, and locations.
Appendix D Annotation Guideline
D.1 Manual QA validation - Selecting Best Answer
The purpose of this task is to evaluate two possible answers for a given question and select the one that is more accurate, clear, and well-structured. Each question will have two answer options, and the task is to choose the better answer based on the evaluation criteria discussed below.
D.1.1 Annotation Task
For each question, there are two answer options. The task is to carefully compare them and select the one that best meets the below criteria:
  1. 1.
    Accuracy. The answer must be factually correct and free from misinformation. It should properly address the question and not introduce unrelated details.
  2. 2.
    Clarity & Readability. The answer should be grammatically correct, well-structured, and easy to read. It should avoid ambiguity, repetition, or fragmented sentences. The information should be presented in a logical and coherent way.
  3. 3.
    Completeness & Relevance. The answer should provide enough relevant information to fully respond to the question. It should avoid unnecessary details that do not contribute to answering the question. If an answer is too short or vague, it may be less useful than a more complete response.
Example
Question: {RLtext}ما الفرق بين برج القاهرة وبرج إيفل؟ Option 1:
{RLtext}برج القاهرة بيتكلم في 100 7 وتمانين متر. اطول من برج ايفل برج ايفل فضل اطول برج في العالم. لمدة 40 سنة طوله كام 300 متر ده 300 5 وتمانين متر. Option 2:
{RLtext}برج القاهرة يبلغ ارتفاعه 187 مترًا، بينما برج إيفل يبلغ ارتفاعه 330 مترًا. برج إيفل كان أطول برج في العالم لمدة 40 عامًا. Evaluation:
Option 2 is the better choice because:
  • It provides accurate numerical information about both towers.
  • It is clear and well-structured, while Option 1 contains fragmented and unclear phrasing.
  • It avoids redundant details and presents the comparison concisely.
Final Selection: Option 2
D.1.2 Annotation Instructions
  • The question and both answer options should be read carefully.
  • Both answers should be evaluated based on the following three main criteria:
    • Accuracy (Is the information correct?)
    • Clarity (Is it well-written and easy to understand?)
    • Completeness & Relevance (Is the question fully and appropriately answered?)
  • The better answer should be selected based on these criteria.
  • If the answers are of equal quality, the one that is more precise and easier to understand should be chosen.
  • If neither answer is acceptable (e.g., both contain major errors), “neither” should be selected, and a comment explaining the reasoning should be provided.
D.1.3 Special Cases and Edge Cases
  • If one answer is factually correct but lacks clarity, while the other is well-written but factually incorrect, the correct answer should be preferred and revised for clarity if necessary.
  • If both answers are equally valid in terms of content, the answer that is clearer and better structured should be selected.
D.2 Manual QA validation - Answer Scoring.
The goal of this task is to assess and, if necessary, revise LLM-edited answers using the information provided in the associated source URLs. Each question includes an LLM-generated answer and its corresponding source. Annotators are expected to verify the quality of the answer based on the source content, and evaluate it according to the criteria outlined below.
Figure A3: An example of the annotation interface for QA annotation for LLM-based QA annotation.
D.2.1 Annotation Task
To assess the quality of LLM-edited answers, we used four different metrics described below. For each metric, assign a score from 1 (very poor) to 5 (excellent), following the descriptions below. Informativeness. Measures the extent to which the assessed text (answer) provides relevant, meaningful, complete (covers all important points), correct (factually accurate), and useful information in response to the given prompt or context. Highly informative content (answer) directly addresses key points; low-informative content (answer) is vague, incomplete, incorrect, or misses critical details.
  • 1 = Not informative: Lacks relevant or meaningful content; does not address the question/context; contains major errors or omissions.
  • 2 = Slightly informative: Provides minimal information; many important details are missing, unclear, or inaccurate.
  • 3 = Moderately informative: Some relevant details are present but lack depth, accuracy, or completeness.
  • 4 = Informative: Well-detailed; addresses most key points clearly, correctly, and thoroughly.
  • 5 = Very informative: Exceptionally detailed, insightful, accurate, and complete; fully addresses all aspects of the prompt/context.
Clarity. Assesses how clearly the assessed text (answer) conveys its meaning. Clear text is well-structured, concise (free from unnecessary information), and easy to understand, free from ambiguity, awkward phrasing, or grammatical errors.
  • 1 = Very unclear: Difficult or impossible to understand; contains major grammatical or structural problems; excessively wordy or redundant.
  • 2 = Somewhat unclear: Some meaning is discernible, but significant ambiguity, awkwardness, or unnecessary verbosity remains.
  • 3 = Neutral: Generally understandable but may require effort or contain minor issues, including some unnecessary length.
  • 4 = Clear: Easy to read and understand with minimal ambiguity; well-structured; mostly concise.
  • 5 = Very clear: Highly readable, precise, concise, and effortlessly understandable.
Plausibility. Refers to whether the answer makes sense and is reasonable in the given context. A plausible response is coherent, logical, correct (factually accurate), and consistent with the input (e.g., question, original text, source image/audio).
  • 1 = Not plausible at all: Response is illogical, unrelated, factually incorrect, or clearly incorrect.
  • 2 = Weakly plausible: Some connection, but contains major logical flaws, inaccuracies, or inconsistencies.
  • 3 = Moderately plausible: Partially logical, but some elements are incomplete, questionable, or partially incorrect.
  • 4 = Plausible: Reasonable, makes sense, factually correct, and is mostly consistent with the context.
  • 5 = Highly plausible: Fully logical, coherent, accurate, and strongly aligned with the context.
Faithfulness. Measures how accurately the assessed text reflects or stays true to the source information or reasoning. Faithful responses avoid adding unrelated details, omitting crucial information (completeness), or misrepresenting facts (correctness and honesty).
  • 1 = Not faithful at all: Largely or entirely unrelated to the source/context; misrepresents information; misses key points or adds incorrect information.
  • 2 = Weakly faithful: Some relevant aspects, but much is misleading, incomplete, or missing.
  • 3 = Moderately faithful: Captures some correct elements, but contains errors, omissions, or partial misrepresentation.
  • 4 = Faithful: Accurately represents the main points and intent of the source/context, without major errors or omissions.
  • 5 = Highly faithful: Fully preserves the original meaning/reasoning, covers all important aspects, and is free from errors or misleading information.
D.3 Annotation Instructions
  • Read carefully: Always review both the context and the text to be assessed as instructed in the platform.
  • Apply each metric independently: Judge informativeness, clarity, plausibility, and faithfulness separately for each item.
  • Be objective: Follow the criteria closely; avoid personal biases or preferences.
  • Consider the context: Assess whether the text meets the needs of the specific task (e.g., Does the answer address the question?).
  • Use full scale: Do not hesitate to assign low or high scores where appropriate. A score of 1 means very poor; 5 means excellent.
D.4 Annotation Interface
Figure A3 present the annotation interface used by annotators to evaluate the edited answers. The source article, shown on the right side of the interface, is used to verify the accuracy and completeness of the answer.
Appendix E Release
The NativQA framework is released under the CC BY-NC-SA 4.0 – Creative Commons Attribution 4.0 International License: https://creativecommons.org/licenses/by-nc-sa/4.0/.